# Quantum Spin Liquids in Pyrochlore Magnets With Non-Kramers Local Moments

### v1

@@ III.4 Generalized Self-Consistent Equations
for every bond $(i,j)$. The approach taken in previous work was to search for a subset of solutions to (25) that respected the PSG constraints [59, 69, 64].
@@ IV.1 Phase Diagram
We compute the mean-field phase diagrams for $\kappa=0.1$ and $J_{zz}^{2}+J_{\pm}^{2}+J_{\pm\pm}^{2}=1$ with $J_{\pm}<0$ and $J_{\pm\pm}>0$, which are presented in Fig. 1 (c) and (d). We chose a small value of $\kappa$ to obtain all the competing QSL phases, since some QSLs may be unstable to magnetic ordering as $\kappa$ increases. We discuss this further in sections IV.2 and V. We also chose $J_{\pm}<0$ to explore the region of the parameter space with frustrated in-plane interactions, where QSLs are more likely to arise. Indeed, numerical studies have shown that the region $J_{\pm}>0$ for $J_{\pm}\geq J_{zz}=0.05$ and $J_{zz}>0$ favours an in-plane pseudospin ferromagnetic order, while the region $J_{\pm}>0$ for $J_{zz}<0$ favours in-plane or Ising-type pseudospin ferromagnetic order [19, 70, 22, 23, 71, 61, 62, 65]. We parameterize the exchange couplings by
@@
Since we have specialized to $n_{1}=0$, there are at most $8$ PSG classes to consider. For the range of parameters we studied, we found that certain PSG classes had identical mean field ansätze in the regime of interest. Namely, the ansätze $(100)$ and $(110)$ were identical as well $(101)$ and $(111)$, reducing the number of PSG classes to $6$.
@@
We point out the phase transition near the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}=-\frac{1}{2}$, $J_{\pm\pm}=0$ where $J_{zz}>0$, which is expected at the classical level [71, 30]. Cluster-variational and pseudo-Majorana fermion functional renormalization group studies of the XXZ model on the pyrochlore lattice also support a phase transition to a magnetically disordered state there as well [61, 62], which supports our finding. Figure 1 (d) shows that the phases $(011)$, $(100)$, and $(001)$ are the second lowest energy phases competing with $(000)$, $(101)$, and $(010)$ respectively. There is also a small region where $(101)$ has the second lowest energy compared to $(000)$. Excluding that small region, the pairs of lowest- and second-lowest-energy PSG classes have the same non-vanishing mean-field parameters $a^{h},\ldots,d^{h},a^{p},\ldots,d^{p}$, defined in Eqs. (18a-18b). However, they can still have distinct physical properties since they have different PSG parameters ${n_{\overline{C}_{6}S}},{n_{ST_{1}}},{n_{\overline{C}_{6}}}$ which govern the spatial structure of the ansätze. Examples of this are shown in the next two sections.
@@ IV.2 Distinguishing Properties of Different PSG Classes
Figure 3: Local frame static spin structure factor components in the $[hhl]$ plane. Rows indicate the PSG class (000), (101), (010), and columns indicate the $xx$, $yy$, and $zz$ components of $\mathcal{S}^{\alpha\beta}_{\textrm{LF}}$ respectively. We set the minimum intensity of each panel to 00 and normalized them independently for better visibility. The inset text shows the normalization factor for each panel relative to its row. For example, in the $(000)$ row, $\mathcal{S}_{\textrm{LF}}^{zz}$ has $0.8$ times the contrast as $\mathcal{S}_{\textrm{LF}}^{xx}$. Momenta are expressed in reciprocal lattice units, and the first Brillouin zone is indicated in black (top left panel).
@@
The spinon dispersions of the six PSG classes are shown in Fig. 2. As discussed in section IV.1, we chose $\kappa=0.1$ so that all QSLs are gapped. However, we note that $(000)$ has a significantly larger bandwidth and smaller energy gap than the other classes, indicating its susceptibility to Schwinger boson condensation. In fact, for $\kappa\approx 0.14$, the class $(000)$ is condensed throughout the entire phase diagram. The minimum of its dispersion occurs at ${{\bm{q}}}=0$, consistent with the ordering wave vector expected from the classical phase diagram [30, 65, 72]. The smaller bandwidth and larger gap of the other PSG classes make them more resistant to condensation, making them more likely to persist at higher values of $\kappa$.
@@
On the other hand, the self-consistent ansätze in the $J_{zz}<0$ region for $(000)$ and $(011)$ have nonzero $\chi,D^{x},D^{y}$, which include $x$ and $y$ triplet mean-field amplitudes. The same is true for $(010)$ and $(001)$. As such, there is no emergent symmetry in these phases, and the spin-spin correlations are different for each component. In contrast to the previous two classes, the intensity of $\mathcal{S}^{zz}_{\textrm{LF}}$ has maximum $[000]$, which is consistent with the ferromagnetic $J_{zz}$ coupling. Figure 3 shows that $\mathcal{S}^{zz}_{\textrm{LF}}$ is significantly stronger than the in-plane correlations for $(010)$ and $(001)$, but weaker than the in-plane correlations for $(000)$ and $(011)$.
@@ IV.3 Dynamic Spin Structure Factor
The observable we are primarily interested in is the dynamic spin structure factor (DSSF) $\mathcal{S}$, which is directly related to the inelastic neutron scattering cross section [28, 73]. The quadrupolar $S^{x}$ and $S^{y}$ do not couple to the neutron’s spin [51], so neutron scattering only probes correlations between $S^{z}$. The resulting DSSF is
@@
We find a high-intensity “rod” pattern (Fig. 4 (a)) for the PSG classes $(000)$ and $(011)$ only when $J_{zz}>0$. They also occur for the PSG classes $(101)$ and $(100)$ for both signs of $J_{zz}$. The high-intensity rods in the $[hhl]$ plane resemble the washed-out pinch point structure observed in experiments [31, 32, 42]. Such rods appear in single tetrahedron calculations of the equal-time structure factor [54]. On the other hand, we find low-intensity rods (Fig. 4 (b)) for $(000)$ and $(011)$ when $J_{zz}<0$, as well as for $(010)$ and $(001)$ regardless of the sign of $J_{zz}$. Since the high-intensity pattern better matches what is experimentally observed, we restrict our attention to the phases $(000),(011)$ (in the region $J_{zz}>0$), and $(101),(100)$ for the rest of this work.
@@
where $\omega_{{\bm{k}}}$ is the spinon dispersion (see Fig. 2). Physically, a neutron that excites two spinons and is scattered with momentum transfer ${{\bm{q}}}$ must transfer an energy $\omega$ that lies within the two-spinon continuum, at least at the mean-field level with non-interacting spinons. The DSSF is therefore 00 outside the two-spinon continuum. The DSSFs of most candidates include a relatively flat feature where most of the spectral weight is concentrated, which is expected from the flatness of the spinon bands as shown in Fig. 2. Interestingly, $(000)$ displays a near-uniform ${{\bm{q}}}$ dependence, at least along the chosen high symmetry path. These DSSFs are broadly consistent with the inelastic signal observed in experiments, but the resolution of currently available instruments may not be fine enough to clearly show any detailed features in Fig. 5.
@@
We also show energy cuts of $\mathcal{S}({{\bm{q}}},\omega)$ integrated over energy windows $(\overline{\omega}-\delta\omega,\overline{\omega}+\delta\omega)$ in the $[hhl]$ plane in Fig. 6. Together, the windows cover the entire two-spinon continuum. Due to the flat features seen in the momentum-resolved DSSF (Fig. 5), the energy window $\overline{\omega}$ centered around the flat feature will have the highest intensity. The cleanest realization of the rod structure occurs in the higher energy windows of $(000)$, but it can also be seen in the other candidates. Some energy-integrated windows display significant modulation of the rods, which is similar to what has been observed in $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ [38].
Figure 6: Energy-integrated DSSF in the $[hhl]$ plane for the lowest and second lowest energy candidates; $(000),(101),(011)$, and $(100)$ in the first, second, third and fourth rows, respectively. Each panel shows $\mathcal{S}({{\bm{q}}},\omega)$ integrated over the energy interval $(\overline{\omega}-\delta\omega,\overline{\omega}+\delta\omega)$ for different $\overline{\omega}$ (inset text) but the same $\delta\omega$ in each row. Together, the $5$ intervals cover the entire two-spinon continuum. We set the minimum intensity of each panel to 00 and normalized them independently for better visibility. Momenta are expressed in reciprocal lattice units, and the first Brillouin zone is shown in white (top left panel).
@@ V Discussion
We investigated possible QSL phases beyond the quantum spin ice regime for pyrochlore oxides with non-Kramers magnetic ions. Local moments are described by pseudospin degrees of freedom consisting of dipolar Ising and quadrupolar XY moments. We used a Schwinger boson mean-field theory and the corresponding projective symmetry group to classify $\mathbb{Z}_{2}$ quantum spin liquids on the pyrochlore lattice with non-Kramers ions. By constructing a mean-field phase diagram, we identified four candidate QSLs in the frustrated region of parameter space that show a high intensity rod structure in the $[hhl]$ plane, consistent with the inelastic signal in neutron scattering experiments [31, 32, 38]. Certain energy-integrated windows of the dynamic spin structure factor show a modulated rod structure that is reminiscent of the features in neutron scattering experiments.
@@
Our phase diagram shows that the $\mathbb{Z}_{2}$ quantum spin liquids $(101)$ and $(100)$ are energetically favoured when $J_{\pm}$ is the dominant coupling. Interestingly, numerical studies of the XXZ model (1) using cluster-variational and pseudo-Majorana fermion functional renormalization group calculations also find a transition to a disordered state past the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}<-\frac{1}{2}$ where $J_{zz}>0$ [61, 62]. This supports the possibility that a quantum spin liquid could occur when $|J_{\pm}|$ becomes comparable to or larger than $|J_{zz}|$. Thus, it is possible that $(101)$ and $(100)$ are among several competing QSL phases in this region of parameter space.
The non-Kramers doublet can be split by crystalline disorder or local strain because its degeneracy is protected only by crystalline symmetries, not by time-reversal symmetry. Previous work indicates that this plays an important role in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ physics and could explain the inelastic neutron scattering signal [33, 42, 56, 57]. It will be important to precisely determine the exchange couplings $J_{zz},J_{\pm},J_{\pm\pm}$ to reliably assess the effects of disorder. At the moment, these parameters are mostly unknown. In the context of Schwinger boson theory, future work could investigate the effects of disorder on candidate QSL phases. On the other hand, our analysis may be more directly applicable to $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$, which is claimed to have minimal disorder [38].
@@ Appendix A Microscopic Model
in terms of standard Mulliken notation [74]. The two-dimensional irreducible representation $E_{g}$ forms the non-Kramers doublet, with an explicit construction given in Eq. (4) [30].
The bond-dependent phase factor $\gamma_{ij}=\gamma_{{{\bm{r}}}_{\mu},{{\bm{r}}}^{\prime}_{\nu}}$ depends only on the sublattice indices $(\mu,\nu)$, and is defined by
@@
$\hat{x}_{\mu}$ $\frac{1}{\sqrt{6}}(-2,1,1)$$\frac{1}{\sqrt{6}}(2,1,1)$ $\frac{1}{\sqrt{6}}(2,1,-1)$ $\frac{1}{\sqrt{6}}(2,-1,1)$
@@ Appendix B Pseudospin Transformations
We show how to derive the matrices $U_{\mathcal{O}}$ for the space group generators and $U_{\mathcal{T}}$ for time-reversal. The transformation of the pseudospin under a space group generator can be deduced by first performing the active transformation on the pseudospin and then changing the local coordinates due to sublattice mixing. The representations of the generators $\overline{C}_{6}$ and $S$ acting on a pseudospin state |±$|\pm\rangle_{\mu}$ on sublattice $\mu$ are
@@
where $\bm{\hat{n}}_{{\mathcal{O}},\mu}$ is the rotation axis of the generator ${\mathcal{O}}$ written in the local frame $\mu$ and $\bm{J}$ is the angular momentum operator. For the ground state (4), we have $J=4$, but we note it is equivalent and simpler computationally to take the ground state doublet as |±$|\pm\rangle=|m=\pm 1\rangle$ with $J=1$ [30]. In global Cartesian coordinates,
@@
Table 3: Relation between the mean-field parameters on the different bonds in the pyrochlore unit cell. The solution of the $\mathbb{Z}_{2}$ PSG is parameterized by $n_{1},n_{C},n_{CS},n_{ST}\in\{0,1\}$, although we set $n_{1}=0$ in this work. Explicit expressions for the matrices $\mathcal{R}^{\gamma}_{ij}$ can be derived as follows: $\mathcal{R}_{ij}^{\gamma}$ is the $4\times 4$ matrix corresponding to the linear transformation of $(a,b,c,d)$ in the row $(i\to j)$, multiplied by the adjacent $\mathbb{Z}_{2}$ phase. For example, $\mathcal{R}_{1_{+},2_{+}}^{h}=\begin{pmatrix}1&&&\\ &-{\frac{1}{2}}&{\frac{\sqrt{3}}{2}}&\\ &{\frac{\sqrt{3}}{2}}&{\frac{1}{2}}&\\ &&&1\end{pmatrix}(-1)^{n_{1}+n_{ST}}$. In this table only, we use that notation $\mu_{\eta}$ to specify sites within the pyrochlore unit cell at $\mathbf{0}$, where $+$ (-) denotes a down- (up-) pointing tetrahedron; see Fig. 1 (a). The mapping back to sublattice indexed pyrochlore coordinates is $\mu_{\eta}\mapsto\mathbf{0}_{\mu}$ if $\eta=+$ and $\mu_{\eta}\mapsto(-{\bm{\hat{e}}}_{\mu})_{\mu}$ if $\eta=-$.
@@
$0_{+}\to 1_{+}$ $(a^{h},b^{h},c^{h},d^{h})$ $1$ $(a^{p},b^{p},c^{p},d^{p})$ $1$
$0_{+}\to 2_{+}$ $(a^{h},-{\frac{1}{2}}b^{h}+{\frac{\sqrt{3}}{2}}c^{h},-{\frac{\sqrt{3}}{2}}b^{h}-{\frac{1}{2}}c^{h},d^{h})$ $1$ $(a^{p},-{\frac{1}{2}}b^{p}+{\frac{\sqrt{3}}{2}}c^{p},-{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ $1$
$0_{+}\to 3_{+}$ $(a^{h},-{\frac{1}{2}}b^{h}-{\frac{\sqrt{3}}{2}}c^{h},{\frac{\sqrt{3}}{2}}b^{h}-{\frac{1}{2}}c^{h},d^{h})$ $1$ $(a^{p},-{\frac{1}{2}}b^{p}-{\frac{\sqrt{3}}{2}}c^{p},{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ $1$
@@
$1_{-}\to 2_{-}$ $(a^{h},-{\frac{1}{2}}b^{h}+{\frac{\sqrt{3}}{2}}c^{h},{\frac{\sqrt{3}}{2}}b^{h}+{\frac{1}{2}}c^{h},-d^{h})$ $(-1)^{n_{1}+n_{ST}}$ $(-a^{p},{\frac{1}{2}}b^{p}-{\frac{\sqrt{3}}{2}}c^{p},-{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ $1$

### v2

@@ III.4 Generalized Self-Consistent Equations
for every bond $(i,j)$. The approach taken in previous work was to search for a subset of solutions to (25) that respected the PSG constraints [Wang2006, Schneider2022, Desrochers2022].
@@ IV.1 Phase Diagram
We compute the mean-field phase diagrams for $\kappa=0.1$ and $J_{zz}^{2}+J_{\pm}^{2}+J_{\pm\pm}^{2}=1$ with $J_{\pm}<0$ and $J_{\pm\pm}>0$, which are presented in Fig. 1 (c) and (d). We chose a small value of $\kappa$ to obtain all the competing QSL phases, since some QSLs may be unstable to magnetic ordering as $\kappa$ increases. We discuss this further in sections IV.2 and V. We also chose $J_{\pm}<0$ to explore the region of the parameter space with frustrated in-plane interactions, where QSLs are more likely to arise. Indeed, numerical studies have shown that the region $J_{\pm}>0$ for $J_{\pm}\geq J_{zz}=0.05$ and $J_{zz}>0$ favours an in-plane pseudospin ferromagnetic order, while the region $J_{\pm}>0$ for $J_{zz}<0$ favours in-plane or Ising-type pseudospin ferromagnetic order [banerjee2008unusual, Kato2015, huang2018dynamics, huang2020extended, Taillefumier2017, benton2018quantum, Schaden2025, Chung2024]. We parameterize the exchange couplings by
@@
Since we have specialized to $n_{1}=0$, there are at most $8$ PSG classes to consider. For the range of parameters we studied, we found that certain PSG classes had identical mean-field ansätze in the regime of interest. Namely, the ansätze $(100)$ and $(110)$ were identical, as well $(101)$ and $(111)$, reducing the number of PSG classes to $6$. Strictly speaking, ansätze belonging to distinct PSG classes are still inequivalent, but we will treat them as the same since the resulting observables are the same. We expect this degeneracy to be lifted upon inclusion of other interactions or next-nearest neighbour terms in the mean-field ansätze.
@@
We point out the phase transition near the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}=-\frac{1}{2}$, $J_{\pm\pm}=0$ where $J_{zz}>0$, which is expected at the classical level [Taillefumier2017, Rau2019]. Cluster-variational and pseudo-Majorana fermion functional renormalization group studies of the XXZ model on the pyrochlore lattice also support a phase transition to a magnetically disordered state there as well [benton2018quantum, Schaden2025], which supports our finding. Figure 1 (d) shows that the phases $(011)$, $(100)$, and $(001)$ are the second lowest energy phases competing with $(000)$, $(101)$, and $(010)$ respectively. There is also a small region where $(101)$ has the second lowest energy compared to $(000)$. Excluding that small region, the pairs of lowest- and second-lowest-energy PSG classes have the same non-vanishing mean-field parameters $a^{h},\ldots,d^{h},a^{p},\ldots,d^{p}$, defined in Eqs. (18a-18b). However, they can still have distinct physical properties since they have different PSG parameters ${n_{\overline{C}_{6}S}},{n_{ST_{1}}},{n_{\overline{C}_{6}}}$ which govern the spatial structure of the ansätze. Examples of this are shown in the next two sections.
@@ IV.2 Distinguishing Properties of Different PSG Classes
Figure 3: Local frame static spin structure factor components in the $[hhl]$ plane. Rows indicate the PSG class (000), (101), (010), and columns indicate the xxxx, yyyy, and $zz$ components of $\mathcal{S}^{\alpha\beta}_{\textrm{LF}}$ respectively. We set the minimum intensity of each panel to 0 and normalized them independently for better visibility. The inset text shows the normalization factor for each panel relative to its row. For example, in the $(000)$ row, $\mathcal{S}_{\textrm{LF}}^{zz}$ has $0.8$ times the contrast as $\mathcal{S}_{\textrm{LF}}^{xx}$. Momenta are expressed in reciprocal lattice units, and the first Brillouin zone is indicated in black (top left panel).
@@
The spinon dispersions of the six PSG classes are shown in Fig. 2. As discussed in section IV.1, we chose $\kappa=0.1$ so that all QSLs are gapped. However, we note that $(000)$ has a significantly larger bandwidth and smaller energy gap than the other classes, indicating its susceptibility to Schwinger boson condensation. In fact, for $\kappa\approx 0.14$, the class $(000)$ is condensed throughout the entire phase diagram. The minimum of its dispersion occurs at ${{\bm{q}}}=0$, consistent with the ordering wave vector expected from the classical phase diagram [Rau2019, Chung2024, yan20017theory]. The smaller bandwidth and larger gap of the other PSG classes make them more resistant to condensation, making them more likely to persist at higher values of $\kappa$.
@@
On the other hand, the self-consistent ansätze in the $J_{zz}<0$ region for $(000)$ and $(011)$ have nonzero $\chi,D^{x},D^{y}$, which include $x$ and $y$ triplet mean-field amplitudes. The same is true for $(010)$ and $(001)$. As such, there is no emergent symmetry in these phases, and the spin-spin correlations are different for each component. In contrast to the previous two classes, the intensity of $\mathcal{S}^{zz}_{\textrm{LF}}$ has a maximum at $[000]$, which is consistent with the ferromagnetic $J_{zz}$ coupling. Figure 3 shows that $\mathcal{S}^{zz}_{\textrm{LF}}$ is significantly stronger than the in-plane correlations for $(010)$ and $(001)$, but weaker than the in-plane correlations for $(000)$ and $(011)$.
@@ IV.3 Dynamic Spin Structure Factor
The observable we are primarily interested in is the dynamic spin structure factor (DSSF) $\mathcal{S}$, which is directly related to the inelastic neutron scattering cross section [Book2011, Boothroyd2020]. The quadrupolar $S^{x}$ and $S^{y}$ do not couple to the neutron’s spin [chen2017dirac], so neutron scattering only probes correlations between $S^{z}$. The resulting DSSF is
@@
We find a high-intensity “rod” pattern (Fig. 4 (a)) for the PSG classes $(000)$ and $(011)$ only when $J_{zz}>0$. They also occur for the PSG classes $(101)$ and $(100)$ for both signs of $J_{zz}$. The high-intensity rods in the $[hhl]$ plane resemble the washed-out pinch point structure observed in experiments [Kimura2013, Petit2016, Wen2017]. Such rods appear in single tetrahedron calculations of the equal-time structure factor [Castelnovo2019]. On the other hand, we find low-intensity rods (Fig. 4 (b)) for $(000)$ and $(011)$ when $J_{zz}<0$, as well as for $(010)$ and $(001)$ regardless of the sign of $J_{zz}$. Since the high-intensity pattern better matches what is experimentally observed, we restrict our attention to the phases $(000),(011)$ (in the region $J_{zz}>0$), and $(101),(100)$ for the rest of this work.
@@
where $\omega_{{\bm{k}}}$ is the spinon dispersion (see Fig. 2). Physically, a neutron that excites two spinons and is scattered with momentum transfer $\bm{q}$}} must transfer an energy $\omega$ that lies within the two-spinon continuum, at least at the mean-field level with non-interacting spinons. The DSSF is therefore 0 outside the two-spinon continuum. The DSSFs of most candidates include a relatively flat feature where most of the spectral weight is concentrated, which is expected from the flatness of the spinon bands as shown in Fig. 2. These DSSFs are broadly consistent with the inelastic signal observed in experiments, but the resolution of currently available instruments may not be fine enough to show any of the detailed features in Fig. 5. Interestingly, the DSSF for $(000)$ has a near-uniform $\bm{q}$}}-dependence, which we attribute to the vanishing of the mean-field amplitudes $D^{x}$ and $D^{y}$ after reaching self-consistency, causing $c^{p}=0$ (as defined in Eq. 18b). In fact, $c^{p}\neq 0$ when $J_{zz}<0$ for $(000)$, and a nontrivial $\bm{q}$}}-dependence in the DSSF is seen.
@@
We also show energy cuts of $\mathcal{S}({{\bm{q}}},\omega)$ integrated over energy windows $(\overline{\omega}-\delta\omega,\overline{\omega}+\delta\omega)$ in the $[hhl]$ plane in Fig. 6. Together, the windows cover the entire two-spinon continuum. Due to the flat features seen in the momentum-resolved DSSF (Fig. 5), the energy window $\overline{\omega}$ centered around the flat feature will have the highest intensity. The cleanest realization of the rod structure occurs in the higher energy windows of $(000)$, but it can also be seen in the other candidates. Some energy-integrated windows display significant modulation of the rods, which is similar to what has been observed in $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$ [Sibille2018].
Figure 6: Energy-integrated DSSF in the $[hhl]$ plane for the lowest and second lowest energy candidates; $(000),(101),(011)$, and $(100)$ in the first, second, third and fourth rows, respectively. Each panel shows $\mathcal{S}({{\bm{q}}},\omega)$ integrated over the energy interval $(\overline{\omega}-\delta\omega,\overline{\omega}+\delta\omega)$ for different $\overline{\omega}$ (inset text) but the same $\delta\omega$ in each row. Together, the 55 intervals cover the entire two-spinon continuum. We set the minimum intensity of each panel to 0 and normalized them independently for better visibility. Momenta are expressed in reciprocal lattice units, and the first Brillouin zone is shown in white (top left panel).
@@ V Discussion
We investigated possible QSL phases beyond the quantum spin ice regime for pyrochlore oxides with non-Kramers magnetic ions. Local moments are described by pseudospin degrees of freedom consisting of dipolar Ising and quadrupolar XY moments. We used a Schwinger boson mean-field theory and the corresponding projective symmetry group to classify $\mathbb{Z}_{2}$ quantum spin liquids on the pyrochlore lattice with non-Kramers ions. By constructing a mean-field phase diagram, we identified four candidate QSLs in the frustrated region of parameter space that show a high intensity rod structure in the $[hhl]$ plane, consistent with the inelastic signal in neutron scattering experiments [Kimura2013, Petit2016, Sibille2018]. Certain energy-integrated windows of the dynamic spin structure factor show a modulated rod structure that is reminiscent of the features in neutron scattering experiments.
@@
Our phase diagram shows that the $\mathbb{Z}_{2}$ quantum spin liquids $(101)$ and $(100)$ are energetically favoured when $J_{\pm}$ is the dominant coupling. Interestingly, numerical studies of the XXZ model (1) using cluster-variational and pseudo-Majorana fermion functional renormalization group calculations also find a transition to a disordered state past the antiferromagnetic Heisenberg point $J_{\pm}/J_{zz}<-\frac{1}{2}$ where $J_{zz}>0$ [benton2018quantum, Schaden2025]. This supports the possibility that a quantum spin liquid could occur when $|J_{\pm}|$ becomes comparable to or larger than $|J_{zz}|$. Thus, it is possible that $(101)$ and $(100)$ are among several competing QSL phases in this region of parameter space.
The non-Kramers doublet can be split by crystalline disorder or local strain because its degeneracy is protected only by crystalline symmetries, not by time-reversal symmetry. Previous work indicates that this plays an important role in $\textrm{Pr}_{2}\textrm{Zr}_{2}\textrm{O}_{7}$ physics and could explain the inelastic neutron scattering signal [Martin2017, Wen2017, Savary2017, Benton2018]. It will be important to precisely determine the exchange couplings $J_{zz},J_{\pm},J_{\pm\pm}$ to reliably assess the effects of disorder. At the moment, these parameters are mostly unknown. In the context of Schwinger boson theory, future work could investigate the effects of disorder on candidate QSL phases. On the other hand, our analysis may be more directly applicable to $\textrm{Pr}_{2}\textrm{Hf}_{2}\textrm{O}_{7}$, which is claimed to have minimal disorder [Sibille2018].
@@ Appendix A Microscopic Model
in terms of standard Mulliken notation [Bradley2009]. The two-dimensional irreducible representation $E_{g}$ forms the non-Kramers doublet, with an explicit construction given in Eq. (4) [Rau2019].
The local basis vectors for the four sublattices of the pyrochlore lattice are shown in table 2. The bond-dependent phase factor $\gamma_{ij}=\gamma_{{{\bm{r}}}_{\mu},{{\bm{r}}}^{\prime}_{\nu}}$ depends only on the sublattice indices $(\mu,\nu)$, and is defined by
@@
$\hat{x}_{\mu}$ $\frac{1}{\sqrt{6}}(-2,1,1)$ -$\frac{1}{\sqrt{6}}(2,1,1)$ $\frac{1}{\sqrt{6}}(2,1,-1)$ $\frac{1}{\sqrt{6}}(2,-1,1)$
@@ Appendix B Pseudospin Transformations
We show how to derive the matrices $U_{\mathcal{O}}$ for the space group generators and $U_{\mathcal{T}}$ for time-reversal. The transformation of the pseudospin under a space group generator can be deduced by first performing the active transformation on the pseudospin and then changing the local coordinates due to sublattice mixing. The representations of the generators $\overline{C}_{6}$ and $S$ acting on a pseudospin state $|\pm\rangle_{\mu}$ on sublattice $\mu$ are
@@
where $\bm{\hat{n}}_{{\mathcal{O}},\mu}$ is the rotation axis of the generator $\mathcal{O}$} written in the local frame $\mu$ and $\bm{J}$ is the angular momentum operator. For the ground state (4), we have $J=4$, but we note it is equivalent and simpler computationally to take the ground state doublet as $|\pm\rangle=|m=\pm 1\rangle$ with $J=1$ [Rau2019]. In global Cartesian coordinates,
@@
Table 3: Relation between the mean-field parameters on the different bonds in the pyrochlore unit cell. The solution of the $\mathbb{Z}_{2}$ PSG is parameterized by $n_{1},n_{C},n_{CS},n_{ST}\in\{0,1\}$, although we set $n_{1}=0$ in this work. Explicit expressions for the matrices $\mathcal{R}^{\gamma}_{ij}$ can be derived as follows: $\mathcal{R}_{ij}^{\gamma}$ is the $4\times 4$ matrix corresponding to the linear transformation of $(a,b,c,d)$ in the row $(i\to j)$, multiplied by the adjacent $\mathbb{Z}_{2}$ phase. For example, $\mathcal{R}_{1_{+},2_{+}}^{h}=\begin{pmatrix}1&&&\\ &-{\frac{1}{2}}&{\frac{\sqrt{3}}{2}}&\\ &{\frac{\sqrt{3}}{2}}&{\frac{1}{2}}&\\ &&&1\end{pmatrix}(-1)^{n_{1}+n_{ST}}$. In this table only, we use that notation $\mu_{\eta}$ to specify sites within the pyrochlore unit cell at $\mathbf{0}$, where ++ (-) denotes a down- (up-) pointing tetrahedron; see Fig. 1 (a). The mapping back to sublattice indexed pyrochlore coordinates is $\mu_{\eta}\mapsto\mathbf{0}_{\mu}$ if $\eta=+$ and $\mu_{\eta}\mapsto(-{\bm{\hat{e}}}_{\mu})_{\mu}$ if $\eta=$-.
@@
$0_{+}\to 1_{+}$ $(a^{h},b^{h},c^{h},d^{h})$ 11 $(a^{p},b^{p},c^{p},d^{p})$ 11
$0_{+}\to 2_{+}$ $(a^{h},-{\frac{1}{2}}b^{h}+{\frac{\sqrt{3}}{2}}c^{h},-{\frac{\sqrt{3}}{2}}b^{h}-{\frac{1}{2}}c^{h},d^{h})$ 11 $(a^{p},-{\frac{1}{2}}b^{p}+{\frac{\sqrt{3}}{2}}c^{p},-{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ 11
$0_{+}\to 3_{+}$ $(a^{h},-{\frac{1}{2}}b^{h}-{\frac{\sqrt{3}}{2}}c^{h},{\frac{\sqrt{3}}{2}}b^{h}-{\frac{1}{2}}c^{h},d^{h})$ 11 $(a^{p},-{\frac{1}{2}}b^{p}-{\frac{\sqrt{3}}{2}}c^{p},{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ 11
@@
$1_{-}\to 2_{-}$ $(a^{h},-{\frac{1}{2}}b^{h}+{\frac{\sqrt{3}}{2}}c^{h},{\frac{\sqrt{3}}{2}}b^{h}+{\frac{1}{2}}c^{h},-d^{h})$ $(-1)^{n_{1}+n_{ST}}$ $(-a^{p},{\frac{1}{2}}b^{p}-{\frac{\sqrt{3}}{2}}c^{p},-{\frac{\sqrt{3}}{2}}b^{p}-{\frac{1}{2}}c^{p},d^{p})$ 11